# Availability-aware Sensor Fusion via Unified Canonical Space for 4D Radar, LiDAR, and Camera

### v1

@@ Abstract
Sensor fusion of camera, LiDAR, and 4-dimensional (4D) Radar has brought a significant performance improvement in autonomous driving (AD). However, there still exist fundamental challenges: deeply coupled fusion methods assume continuous sensor availability, making them vulnerable to sensor degradation and failure, whereas sensor-wise cross-attention fusion methods struggle with computational cost and unified feature representation. This paper presents availability-aware sensor fusion (ASF), a novel method that employs unified canonical projection (UCP) to enable consistency in all sensor features for fusion and cross-attention across sensors along patches (CASAP) to enhance robustness of sensor fusion against sensor degradation and failure. As a result, the proposed ASF shows a superior object detection performance to the existing state-of-the-art fusion methods under various weather and sensor degradation (or failure) conditions; Extensive experiments on the K-Radar dataset demonstrate that ASF achieves improvements of 9.7% in $AP_{BEV}$ (87.2%) and 20.1% in $AP_{3D}$ (73.6%) in object detection at IoU=0.5, while requiring a low computational cost. The code will be available at https://github.com/kaist-avelab/K-Radar.
Figure 1: Visualization of feature representation with t-SNE [1] at different stages of ASF for ‘Sedan’ class. Red, green, blue, and gray dots represent features from camera, LiDAR, 4D Radar, and fused features, respectively. Symbols in solid lines such as circle and triangle, square, and star indicate normal, sleet, and heavy snow conditions, respectively. (a) Initial output features from sensor-specific encoders show inconsistent distribution across sensors. (b) After UCP, features become better aligned to the fused feature. (c) After CASAP-PN, features from available sensors form cohesive clusters (in dashed symbols) based on weather conditions. Note that in adverse weather condition, camera features show larger deviation due to the degradation. Additional t-SNE visualizations are in Appendix B.
Figure 2: Comparison of sensor fusion methods: (a) DCF (e.g., 3D-LRF [2]), (b) SCF (e.g., DPFT [3]), and (c) ASF. FV, Obj., GT, $\mathcal{L}_{cls}$ and $\mathcal{L}_{reg}$ stand for ‘front-view’, ‘objects’, ‘ground truths’, ‘classification loss’, and ‘regression loss’, respectively. ‘Feature coupling’ refers to methods that combine features from multiple sensors to create new features. Optional components are in dashed lines; for example, [4] combines camera and LiDAR features without transforming the camera viewpoint, while [5] fuses features through a transformer decoder head [6] without explicit feature coupling. ASF does not apply feature coupling to ensure independence between sensors.
@@ 1 Introduction
Autonomous driving (AD) technology has advanced rapidly, with multiple companies adopting multi-sensor fusion approaches that combine two or more sensors, such as cameras, LiDAR, and 4-dimensional (4D) Radar, to achieve more robust and reliable perception [7, 8, 9]. Cameras uniquely provide color information but struggle with depth estimation; LiDAR delivers high-resolution 3-dimensional (3D) point cloud data but is less reliable in adverse weather conditions [10]; and 4D Radar, despite its relatively low angular resolution, offers robustness in adverse weather and directly measures relative velocity [11, 12, 13, 14, 15, 16]. This complementary nature initiated sensor fusion between camera, LiDAR, and 4D Radar to improve perception performance and reliability compared to a single-sensor [5, 17, 18, 19, 20].
@@
Most multi-modal sensor fusion methods can be categorized into two methods. The first is deeply coupled fusion (DCF), which directly combines feature maps (FMs) extracted by sensor-specific encoder, as illustrated in Fig. 2-(a). While it is simple to implement and computationally efficient with excellent performance in various benchmarks [2, 18, 17, 21, 22, 23], it assumes all sensors are functioning properly and consistently. This makes DCF vulnerable to sensor degradation due to adverse weathers, surface-damages, and sensor failure. Moreover, DCF requires retraining the entire neural network when the number of sensors changes, as the size of the fused FM (i.e., the input to the detection head) changes. The second method is sensor-wise cross-attention fusion (SCF), which divides features extracted from each sensor into patches with positional-embedding (e.g., depth information for camera [24, 25]) and selectively combines available patches using cross-attention (Fig. 2-(b)), allowing it to handle cases where some sensors are degraded. However, SCF does not have sensor scalability [26], since the method does not project sensor-specific features into a standardized representation [27, 28, 29, 30, 31], and SCF incurs computational complexity that increases quadratically with the number of patches [3, 5, 32].
@@
One of the fundamental limitations of existing fusion methods stems from inconsistencies in feature representation across different sensors [26, 33]. Cameras produce 2D RGB images, whereas LiDAR generates 3D point clouds, and 4D Radar produces low-resolution tensors with power values. Therefore, each sensor extracts features with different representations for the same object, making consistent fusion challenging (as shown in Fig. 1-(a)). To address this inconsistency, an ideal strategy could project features from different sensors into a unified canonical space for fusion. The concept of ‘True Redundancy [26]’, that ensures sensor independence while maintaining canonical feature representation for any sensor combination, suggests a promising direction for highly reliable and robust sensor fusion.
@@
Therefore, we propose availability-aware sensor fusion (ASF) method (Fig. 2-(c)), in which each sensor (i.e., camera, Lidar, and 4D Radar) performs independently while being complementarily fused through a projection to a unified canonical space. As a result, the proposed method addresses the limitations of both DCF and SCF simultaneously. The key innovation of ASF is in two sub-modules; First, unified canonical projection (UCP) projects features from each sensor into a unified space based on common criteria (i.e., canonical). Since UCP is optimized using the same reference query for all sensors, inconsistencies in sensor features are eliminated. While Fig. 1-(a) shows sensor features represented without clear patterns, Fig. 1-(b) demonstrates how UCP aligns the features from each sensor to the fused feature. Second, cross-attention across sensors along patches with post-feature normalization (CASAP-PN) estimates the availability of sensors through patch-wise cross-attention on features projected into the unified canonical space, assigning higher weights to features from available sensors and lower weights to features from missing or uncertain sensors. Unlike SCF that applies the cross-attention across all sensors ($N_{s}$) and patches ($N_{p}$) (i.e., $O(N_{s}^{2}N_{p}^{2})$), ASF only applies the cross-attention across sensors along patches (i.e., $O(N_{s}^{2}N_{p})$). Because of this, ASF eliminates complex positional-embedding and improves computationally efficiency. Additionally, it applies normalization to ensure that (camera, LiDAR, and 4D Radar) features can be processed consistently by the detection head regardless of sensor combination. In contrast to Fig. 1-(b), where each sensor’s features remain separate, Fig. 1-(c) shows how sensor features cluster together after post-feature normalization (except for camera images become useless in adverse weather conditions, as shown in Fig. 3). This enables ASF to flexibly handle sensor degradation or failure and to achieve the reliability embodied in the ‘True Redundancy’ concept [26] for AD implementation.
@@
To integrate the availability awareness into the detection head, we propose a sensor combination loss (SCL) that optimizes learning outcomes across all sensor combinations. SCL considers individual sensor unavailability during the training, enabling the system to maintain high performance in the presence of unexpected sensor failure or adverse weather conditions. The effectiveness of our proposed ASF method has been validated on the K-Radar dataset [12], demonstrating improvements of 9.7% in $AP_{BEV}$ (87.2%) and 20.1% in $AP_{3D}$ (73.6%) for detection performance at IoU=0.5 compared to state-of-the-art (SOTA) fusion methods [2, 34], which includes the performance in extreme situations such as sensor degradation or failure (unavailable).
The main contributions of this paper are as follows:
@@
  • We propose ASF based on UCP and CASAP-PN that achieves superior performance to SOTA methods and robust performance against sensor degradation and failure.
  • We propose SCL for the loss function to optimize the detection performance for all possible sensor combinations.
  • Through extensive experiments on the K-Radar dataset, we demonstrate that ASF achieves the high performance with low computational load.
@@
The remainder of this paper is organized as follows: Section 2 introduces existing methods, and Section 3 describes the proposed ASF in detail. Section 4 presents experimental settings and results on the K-Radar dataset, analyzing performance in various scenarios including sensor degradation and failure. Section 5 concludes the paper with a summary and discusses the limitations. All code and logs for ASF will be made available at the K-Radar github repository.
@@ 2.1 Deeply Coupled Fusion (DCF)
DCF constructs a fused feature map (FM) by concatenating FMs from each sensor. Most studies focus on fusing camera with LiDAR or 4D Radar [18, 17, 19, 20], or combining LiDAR and 4D Radar [2, 34]. Implementations range from directly fusing the front-view (FV) camera image with LiDAR or 4D Radar without view transformation [4] to applying learnable BEV transforms [35] and concatenating at the BEV stage [18, 17, 19]. DCF method is straightforward to implement and computationally efficient compared to SCF, demonstrating strong performance across multiple benchmarks.
Recent DCF studies have improved performance by applying feature coupling, where the FMs of each sensor are enhanced with FMs of other sensors using multilayer perception (MLP) or attention mechanisms [36]. 3D-LRF [2] demonstrated superior performance to the conventional DCF methods [37, 17, 18] by applying attention between LiDAR and 4D Radar voxel features before concatenation. L4DR [34] achieved SOTA performance on the K-Radar dataset by weighting each sensor’s FM using coupled BEV FMs with LiDAR and 4D Radar. However, DCF assumes that all sensors are functioning normally, as they rely on the fused FM constructed by concatenating FMs from all sensors, which makes DCF performance vulnerable to sensor degradation and failure.
@@ 2.2 Sensor-wise Cross-attention Fusion (SCF)
SCF divides FMs of all sensors into patches and dynamically combines patches of all sensors through cross-attention in a transformer decoder head [3, 32, 5]. This method can inherently utilize sensor availability through the cross-attention. TransFusion [32] is the first SCF method that addresses sensor availability, but its sequential fusion method (e.g., performing LiDAR detection first and then fusing with camera data) makes inference impossible when LiDAR is unavailable. CMT [5] presents an availability-aware sensor fusion of camera and LiDAR data using a transformer decoder head without applying feature coupling to individual sensors. However, without feature coupling for camera, CMT relies on positional embedding to incorporate depth information into camera FMs [24], which results in 3D patches ($H\!\times\!W\!\times\!D$). This leads to computational complexity of $O(N_{s}^{2}N_{p}^{2})$ (where $N_{s}$ and $N_{p}$ are the number of sensors and patches, respectively), causing explosive growth in computational cost and memory usage. For instance, CMT requires 8 A100 GPUs with 80GB VRAM to train with a batch size of 16. Recently, DPFT [3] projects depth-aware 4D Radar FMs onto the FM from the front-view camera image to create patches and verify sensor availability. However, it achieves relatively low performance (56.1% $AP_{3D}$ at IoU=0.3) but demands high computational cost that scales quadratically with the number of patches.
In summary, SCF methods require high computational cost and directly utilize FM from each sensor without a projection process, making it difficult to create canonical FMs in a unified space and to improve sensor scalability. This makes SCF methods challenging to implement in resource-constrained real-time AD systems.
@@ 3.1 Sensor Fusion Framework
Focusing on our ASF contribution, we utilize established methods for the sensor-specific encoders and detection head. Specifically, we adopt BEVDepth [38], SECOND [39], and RTNH [12] backbones for camera, LiDAR, and 4D Radar, respectively, along with a SSD detection head [40, 41]. Further specifications regarding the overall sensor fusion framework, such as sensor-specific encoders and detection head, are provided in Appendix A.
@@ 3.2 Availability-aware Sensor Fusion (ASF)
As illustrated in Fig. 2-(c), the proposed ASF consists of two key components: unified canonical projection (UCP) and cross-attention across sensors along patches with post-feature normalization (CASAP-PN).
Unified Canonical Projection (UCP). One of the key challenge in multi-modal sensor fusion is the inherent inconsistency of features from different sensors [26, 33], as visualized in Fig. 1-(a). To tackle this, we divide BEV FMs into an equal number of patches for all sensors and train projection functions to transform features from each sensor into a unified space based on the same criteria (i.e., reference query in CASAP-PN). To formally define our methods, we first represent the same-sized FMs of each sensor as:
@@
where $C_{s}$ denotes the channel dimension for sensor $s$, $H$ and $W$ represent the identical height and width of the BEV FMs for all sensors, respectively, and $S_{C}$, $S_{L}$, and $S_{R}$ denote camera, LiDAR, and 4D Radar sensors, respectively. Each FM is then divided into patches $\mathbf{F}_{p}^{s}$ with height $P_{H}$ and width $P_{W}$:
@@
where $\mathcal{T}_{p}(\cdot)$ is the operation that divides each FM into patches, $N_{p}\!=\!(H/P_{H})\!\times\!(W/P_{W})$ is the number of patches, which is identical across all sensors since each FM has the same spatial size, and ‘$1\!:\!N_{p}$’ denotes ‘$1,2,...,N_{p}$’. Note that since the patches are already spatially aligned (i.e., $\mathbf{F}_{p,i}^{S_{C}}$, $\mathbf{F}_{p,i}^{S_{L}}$, and $\mathbf{F}_{p,i}^{S_{R}}$ correspond to the same position), our method eliminates the use of computationally expensive positional-embedding [24] required for SCF [5, 3]. Then, we apply a parallel operation along patches that projects each patch to have the same channel dimension $C_{u}$. This is the UCP operation $\mathcal{U}^{s}(\cdot)$ that transforms sensor-specific patches into patches in a unified canonical space. The UCP-processed patch $\mathcal{P}_{u}^{s}$ for each sensor is expressed as follows:
@@
where LN and $n_{u}$ denote layer normalization [42] for training stability and the number of sequential projection functions incorporating MLP for transformation and GeLU[43] for non-linearity, respectively. While our framework allows for repetition of the projection function to increase non-linearity, with 1 or 2 repetitions being sufficient (we use $n_{u}=2$, which aligns features as demonstrated in Fig. 1-(b)). Note that $\mathcal{U}^{s}$ is trained separately for each sensor based on reference query, which results in alignment of features from all sensors with respect to the fused feature as shown in Fig. 1-(b).
Cross-attention Across Sensors Along Patches with Post-feature Normalization (CASAP-PN). The patches $\mathbf{F}_{u}^{s}$ projected into the unified canonical space by UCP serve as keys ($K$) and values ($V$) for a trainable reference query $\mathbf{Q}_{ref}\!\!\in\!\mathbb{R}^{N_{q}\times C_{u}}$ (where $N_{q}$ is the number of queries $Q$), and we perform cross-attention across sensors along patches as:
$\mathbf{Q}_{ref,i}^{\prime}\!=\!\mathrm{CrossAttn}($ $Q\!=\!\mathbf{Q}_{ref},K\&V\!\in\!$ (5)
$\{\mathbf{F}_{u,i}^{S_{C}},\mathbf{F}_{u,i}^{S_{L}},\mathbf{F}_{u,i}^{S_{R}}\}),i\!=\!1:N_{p},$
@@
Compared to ASF, cross-attention in SCF is performed across all positional-embedded patches with respect to object queries $\mathbf{Q}_{obj}$ in the transformer decoder head. This can be mathematically expressed as:
$\mathbf{Q}_{obj}^{\prime}\!=\!\mathrm{CrossAttn}^{(n_{td})}(Q\!=\!\mathbf{Q}_{obj},K\&V\!\in\!$ (6)
$\{\mathbf{F}_{pe,1}^{S_{C}}\!,\!...,\!\mathbf{F}_{pe,N_{p}}^{S_{C}}\!,\!\mathbf{F}_{pe,1}^{S_{L}}\!,\!...,\!\mathbf{F}_{pe,N_{p}}^{S_{L}}\!,\!\mathbf{F}_{pe,1}^{S_{R}}\!,\!...,\!\mathbf{F}_{pe,N_{p}}^{S_{R}}\!\}\!),$
where $\mathbf{F}_{pe}^{s}$ represents patches with positional-embedding, and $n_{td}$ (usually larger than 6 [24, 5]) is the number of stacked transformer decoders. Eq. 6 shows the cross-attention across all sensors and all patches (i.e., the K&V set contains $N_{s}N_{p}$ patches, resulting in $O(N_{s}^{2}N_{p}^{2})$ computational complexity). In contrast, in Eq. 5, the cross-attention is applied across sensors and along patches, which requires only $O(N_{s}^{2}N_{p})$ computational operations. This is a significant computational costs reduction as $N_{s}\ll N_{p}$. Moreover, as demonstrated in Fig. 3 and Tab. 1, ASF achieves better performance with only a single cross-attention layer than SCF utilizing stacked cross-attention layers (i.e., $n_{td}\!\geq\!6$).

### v2

@@ Abstract
Sensor fusion of camera, LiDAR, and 4-dimensional (4D) Radar has brought a significant performance improvement in autonomous driving. However, there still exist fundamental challenges: deeply coupled fusion methods assume continuous sensor availability, making them vulnerable to sensor degradation and failure, whereas sensor-wise cross-attention fusion methods struggle with computational cost and unified feature representation. This paper presents availability-aware sensor fusion (ASF), a novel method that employs unified canonical projection (UCP) to enable consistency in all sensor features for fusion and cross-attention across sensors along patches (CASAP) to enhance robustness of sensor fusion against sensor degradation and failure. As a result, the proposed ASF shows a superior object detection performance to the existing state-of-the-art fusion methods under various weather and sensor degradation (or failure) conditions. Extensive experiments on the K-Radar dataset demonstrate that ASF achieves improvements of 9.7% in $AP_{BEV}$ (87.2%) and 20.1% in $AP_{3D}$ (73.6%) in object detection at IoU=0.5, while requiring a low computational cost. All codes are available at https://github.com/kaist-avelab/k-radar.
@@ 1 Introduction
Autonomous driving technology has advanced rapidly, with multiple companies adopting multi-sensor fusion approaches that combine two or more sensors, such as cameras, LiDAR, and 4-dimensional (4D) Radar, to achieve more robust and reliable perception (self_driving_survey, sensor_fusion_survey). Cameras uniquely provide color information but struggle with depth estimation; LiDAR delivers high-resolution 3-dimensional (3D) point cloud data but is less reliable in adverse weather conditions (adverse_weahter_survey); and 4D Radar, despite its relatively low angular resolution, offers robustness in adverse weather and directly measures relative velocity (kradar, vod, rtnhp, survey_4d_radar). This complementary nature initiated sensor fusion between camera, LiDAR, and 4D Radar to improve perception performance and reliability compared to a single-sensor (cmt, bevfusion_nips, bevfusion_icra, rcfusion).
@@
Most multi-modal sensor fusion methods can be categorized into two methods. The first is deeply coupled fusion (DCF), which directly combines feature maps (FMs) extracted by sensor-specific encoder, as illustrated in Fig. 1-(a). While it is simple to implement and computationally efficient with excellent performance in various benchmarks (lrf, bevfusion_icra, bevfusion_nips, simplebev, nuscenes, kitti), it assumes all sensors are functioning properly and consistently. This makes DCF vulnerable to sensor degradation due to adverse weathers, surface-damages, and sensor failure. Moreover, DCF requires retraining the entire neural network when the number of sensors changes, as the size of the fused FM (i.e., the input to the detection head) changes. The second method is sensor-wise cross-attention fusion (SCF), which divides features extracted from each sensor into patches with positional-embedding (e.g., depth information for camera (petr, detr3d)) and selectively combines available patches using cross-attention (Fig. 1-(b)), allowing it to handle cases where some sensors are degraded. However, SCF does not have sensor scalability, since the method does not project sensor-specific features into a standardized representation (unified_segmentation, unified_vision_language, unified_image_text). In addition, SCF incurs computational complexity that scales with the number of patches, resulting in substantial computational overhead when processing multiple sensors with numerous patches (dpft, cmt, transfusion).
@@
One of the fundamental limitations of existing fusion methods stems from inconsistencies in feature representation across different sensors (mobileye, sensor_fusion_survey_inconsistence). Cameras produce 2D RGB images, whereas LiDAR generates 3D point clouds, and 4D Radar produces low-resolution tensors with power values. Therefore, each sensor extracts features with different representations for the same object, making consistent fusion challenging (as shown in Fig. 2-(a)). To address this inconsistency, an ideal strategy could project features from different sensors into a unified canonical space for fusion. The concept of ‘True Redundancy (mobileye)’, that ensures sensor independence while maintaining canonical feature representation for any sensor combination, suggests a promising direction for highly reliable and robust sensor fusion.
Figure 1: Comparison of sensor fusion methods: (a) DCF (e.g., 3D-LRF (lrf)), (b) SCF (e.g., CMT (cmt)), and (c) ASF. FV, BEV, Obj., GT, $\mathcal{L}_{cls}$ and $\mathcal{L}_{reg}$ stand for ‘front-view’, ‘bird’s eye-view’, ‘objects’, ‘ground truths’, ‘classification loss’, and ‘regression loss’, respectively. ‘Feature coupling’ refers to methods that combine features from multiple sensors to create new features. Optional components are in dashed lines; for example, (pointpainting) combines camera and LiDAR features without transforming the camera viewpoint, while (cmt) fuses features through a transformer decoder head (detr) without explicit feature coupling. ASF does not apply feature coupling to ensure independence between sensors.
Figure 2: Visualization of feature representation with t-SNE (tsne) at different stages of ASF for ‘Sedan’ class. Red, green, blue, and gray dots represent features from camera, LiDAR, 4D Radar, and fused features, respectively. Symbols in solid lines such as circle and triangle, square, and star indicate normal, sleet, and heavy snow conditions, respectively. (a) Initial output features from sensor-specific encoders show inconsistent distribution across sensors. (b) After unified canonical projection (UCP), features become better aligned to the fused feature. (c) After cross-attention across sensors along patches (CASAP), features from available sensors form cohesive clusters (in dashed symbols) based on weather conditions. Note that in adverse weather, camera features show larger deviation due to the degradation. Additional visualizations are in Appendix B.
@@
Therefore, we propose availability-aware sensor fusion (ASF) method (Fig. 1-(c)), in which each sensor performs independently while being complementarily fused through a projection to a unified canonical space. As a result, the proposed method addresses the limitations of both DCF and SCF simultaneously. The key innovation of ASF is in two sub-modules; First, unified canonical projection (UCP) projects features from each sensor into a unified space based on common criteria (i.e., canonical). Since UCP is optimized using the same reference query for all sensors, inconsistencies in sensor features are eliminated. While Fig. 2-(a) shows sensor features represented without clear patterns, Fig. 2-(b) demonstrates how UCP aligns the features from each sensor to the fused feature. Second, cross-attention across sensors along patches (CASAP) estimates the availability of sensors through patch-wise cross-attention on features projected into the unified canonical space, assigning higher weights to features from available sensors and lower weights to features from missing or degraded sensors. Unlike SCF that applies the cross-attention across all sensors ($N_{s}$) and patches ($N_{p}$) (i.e., $O(N_{q}N_{s}N_{p})$ for $N_{q}$ queries), ASF only applies the cross-attention across sensors along patches (i.e., $O(N_{q}N_{s})$). Because of this, ASF eliminates complex positional-embedding and improves computationally efficiency. Additionally, it applies normalization to ensure that (camera, LiDAR, and 4D Radar) features can be processed consistently by the detection head regardless of sensor combination. In contrast to Fig. 2-(b), where each sensor’s features remain separate, Fig. 2-(c) shows how sensor features cluster together after CASAP (except for camera become useless in adverse weathers, as shown in Fig. 3). This enables ASF to flexibly handle sensor degradation or failure and to achieve the reliability embodied in the ‘True Redundancy’ concept (mobileye) for autonomous driving implementation.
@@
To integrate the availability awareness into the detection head, we propose a sensor combination loss (SCL) that optimizes learning outcomes across all sensor combinations. SCL considers individual sensor unavailability during the training, enabling the system to maintain high performance in the presence of unexpected sensor failure or adverse weather conditions. The effectiveness of our proposed ASF method has been validated on the K-Radar dataset (kradar), demonstrating improvements of 9.7% in $AP_{BEV}$ (87.2%) and 20.1% in $AP_{3D}$ (73.6%) for detection performance at IoU=0.5 compared to state-of-the-art (SOTA) methods (lrf, l4dr), which includes the performance in extreme situations such as sensor degradation or failure (i.e., unavailable).
@@
The main contributions of this paper are as follows: (1) We propose ASF based on UCP and CASAP that achieves superior performance to SOTA methods and robust performance against sensor degradation and failure. (2) We propose SCL for the loss function to optimize the detection performance for all possible sensor combinations. (3) Through extensive experiments on the K-Radar dataset, we demonstrate that ASF achieves the high performance with low computational load.
@@
The remainder of this paper is organized as follows: Section 2 introduces existing methods, and Section 3 describes the proposed ASF in detail. Section 4 presents experimental settings and results on the K-Radar dataset, analyzing performance in various scenarios including sensor degradation and failure. Section 5 concludes the paper with a summary and discusses the limitations. All codes and logs for ASF are available at https://github.com/kaist-avelab/k-radar.
@@ 2.1 Deeply Coupled Fusion (DCF)
DCF constructs a fused feature map (FM) by concatenating FMs from each sensor. Most studies focus on fusing camera with LiDAR or 4D Radar (bevfusion_icra, bevfusion_nips, rcfusion, lxl), or combining LiDAR and 4D Radar (lrf, l4dr). Implementations range from directly fusing the front-view (FV) camera image with LiDAR or 4D Radar without view transformation (pointpainting) to applying learnable BEV transforms (lss) and concatenating at the BEV stage (bevfusion_icra, bevfusion_nips, rcfusion). DCF method is straightforward to implement and computationally efficient compared to SCF, demonstrating strong performance across multiple benchmarks.
Recent DCF studies have improved performance by applying feature coupling, where the FMs of each sensor are enhanced with FMs of other sensors using multi-layer perception (MLP) or attention mechanisms (attention). 3D-LRF (lrf) demonstrated superior performance to the conventional DCFs (bevfusion_nips, bevfusion_icra) by applying attention between LiDAR and 4D Radar voxel features before concatenation. L4DR (l4dr) achieved SOTA performance on the K-Radar dataset by weighting each sensor’s FM using coupled BEV FMs with LiDAR and 4D Radar. However, DCF assumes all sensors are functioning normally, as they rely on the fused FM constructed by concatenating FMs from all sensors, making DCF vulnerable to sensor degradation and failure. This limitation arises because the training process does not expose the model to potential sensor degradation or failure scenarios that commonly occur during deployment.
@@ 2.2 Sensor-wise Cross-attention Fusion (SCF)
SCF divides sensor-specific FMs into patches and dynamically combines them through cross-attention in a transformer decoder head (cmt), inherently accommodating sensor availability. TransFusion (transfusion) is the first SCF that addresses sensor availability, but its sequential fusion method (e.g., performing LiDAR detection first and then fusing with camera data) makes inference impossible when LiDAR is unavailable. CMT (cmt) presents an availability-aware sensor fusion of camera and LiDAR data using a transformer decoder head without applying feature coupling to individual sensors. However, without feature coupling for camera, CMT relies on positional embedding to incorporate depth information into camera FMs (petr), which results in 3D patches ($H\!\times\!W\!\times\!D$). This leads to computational complexity of $O(N_{q}N_{s}N_{p})$ (where $N_{q}$, $N_{s}$, and $N_{p}$ are the number of queries, sensors, and patches, respectively), causing explosive growth in computational cost and memory usage. For instance, CMT requires 8 A100 GPUs with 80GB VRAM to train with a batch size of 16.
Recently, DPFT (dpft) creates independent FMs and projects sensor-agnostic query points onto different FMs to verify sensor availability, achieving 56.1% $AP_{3D}$ at IoU=0.3 (using only 4D Radar and camera). Unlike methods that utilize entire FMs, DPFT achieves reasonable computational efficiency by employing deformable attention (deformable_attention) that considers varying receptive fields using only a small number of key points. However, similar to CMT, DPFT performs object detection using variable object queries, which does not establish a common representation across different sensors (as illustrated in Fig. 2).
@@ 3.1 Sensor Fusion Framework
Focusing on our ASF contribution, we utilize established methods for the sensor-specific encoders and detection head. Specifically, we adopt BEVDepth (bevdepth), SECOND (second), and RTNH (kradar) backbones for camera, LiDAR, and 4D Radar, respectively, along with a SSD detection head (ssd). Further specifications regarding the overall sensor fusion framework, such as sensor-specific encoders and detection head, are provided in Appendix A.
@@ 3.2 Availability-aware Sensor Fusion (ASF)
As illustrated in Fig. 1-(c), the proposed ASF consists of two key components: unified canonical projection (UCP) and cross-attention across sensors along patches (CASAP).
Unified Canonical Projection (UCP). One of the key challenge in multi-modal sensor fusion is the inherent inconsistency of features from different sensors (mobileye, sensor_fusion_survey_inconsistence), as visualized in Fig. 2-(a). To tackle this, we divide BEV FMs into an equal number of patches for all sensors and train projection functions to transform features from each sensor into a unified space based on the same criteria (i.e., reference query in CASAP). To formally define our methods, we first represent the same-sized FMs of each sensor as:
@@
where $C_{s}$ denotes the channel dimension for sensor $s$, $H$ and $W$ represent the identical height and width of BEV FMs for all sensors, respectively, and $S_{C}$, $S_{L}$, and $S_{R}$ denote camera, LiDAR, and 4D Radar sensors, respectively. Each FM is then divided into patches $\mathbf{F}_{p}^{s}$ with height $P_{H}$ and width $P_{W}$:
@@
where $\mathcal{T}_{p}(\cdot)$ is the operation that divides each FM into patches, $N_{p}\!=\!(H/P_{H})\!\times\!(W/P_{W})$ is the number of patches, which is identical across all sensors since each FM has the same spatial size, and $1\!:\!N_{p}$’ denotes $1,2,...,N_{p}$’. Note that since the patches are already spatially aligned (i.e., $\mathbf{F}_{p,i}^{S_{C}}$, $\mathbf{F}_{p,i}^{S_{L}}$, and $\mathbf{F}_{p,i}^{S_{R}}$ correspond to the same position), our method eliminates the use of computationally expensive positional-embedding (petr) required for SCF (cmt, dpft). Then, we apply a parallel operation along patches that projects each patch to have the same channel dimension $C_{u}$. This is the UCP operation $\mathcal{U}^{s}(\cdot)$ that transforms sensor-specific patches into patches in a unified canonical space. The UCP-processed patch $\mathcal{P}_{u}^{s}$ for each sensor is expressed as:
@@
where LN and $n_{u}$ denote layer normalization (layernorm) for training stability and the number of sequential projection functions incorporating MLP for transformation and GeLU (gelu) for non-linearity, respectively. While our framework allows for repetition of the projection function to increase non-linearity, with 1 or 2 repetitions being sufficient (we use $n_{u}=2$, which aligns features as demonstrated in Fig. 2-(b)). Note that $\mathcal{U}^{s}$ is trained separately for each sensor based on reference query, which results in alignment of features from all sensors with respect to the fused feature as shown in Fig. 2-(b).
Cross-attention Across Sensors Along Patches (CASAP). The patches $\mathbf{F}_{u}^{s}$ projected into the unified canonical space by UCP serve as keys ($K$) and values ($V$) for a trainable reference query $\mathbf{Q}_{ref}\!\!\in\!\mathbb{R}^{N_{q}\times C_{u}}$ (where $N_{q}$ is the number of queries $Q$), and we perform cross-attention across sensors along patches as:
$\mathbf{Q}_{ref,i}^{\prime}\!=\!\mathrm{CrossAttn}(Q\!=\!\mathbf{Q}_{ref},K\&V\!\in\!\{\mathbf{F}_{u,i}^{S_{C}},\mathbf{F}_{u,i}^{S_{L}},\mathbf{F}_{u,i}^{S_{R}}\}),i\!=\!1:N_{p},$ (5)
@@
Compared to ASF, cross-attention in SCF is performed across all patches with respect to object queries $\mathbf{Q}_{obj}$ in the transformer decoder head. This can be mathematically expressed as:
$\mathbf{Q}_{obj}^{\prime}\!\!=\!\!\mathrm{CrossAttn}^{(n_{td})}\!(Q\!=\!\mathbf{Q}_{obj},K\&V\!\!\in\!\!\{\mathbf{F}_{pe,1}^{S_{C}}\!,\!...,\!\mathbf{F}_{pe,N_{p}}^{S_{C}}\!,\!\mathbf{F}_{pe,1}^{S_{L}}\!,\!...,\!\mathbf{F}_{pe,N_{p}}^{S_{L}}\!,\!\mathbf{F}_{pe,1}^{S_{R}}\!,\!...,\!\mathbf{F}_{pe,N_{p}}^{S_{R}}\!\}\!),$ (6)
where $\mathbf{F}_{pe}^{s}$ represents patches with positional-embedding, and $n_{td}$ (usually larger than 6 (petr, cmt)) is the number of stacked transformer decoders. Eq. 6 shows the cross-attention across all sensors and all patches (i.e., the K&VK\&V set contains $N_{s}N_{p}$ patches, resulting in $O(N_{q}N_{s}N_{p})$ computational complexity for $N_{q}$ queries). In contrast, in Eq. 5, the cross-attention is applied across sensors and along patches, which requires only $O(N_{q}N_{s})$ computational operations. This is a significant computational costs reduction as $N_{s}\ll N_{p}$. Moreover, as demonstrated in Fig. 3 and Tab. 1, ASF achieves better performance with only a single cross-attention layer than SCF utilizing stacked cross-attention layers (i.e., $n_{td}\!\geq\!6$).